OpenStudy (anonymous):

Find the vertex of f(x)=-2x^2+2x-3 Find the Vertex (h,k) h= a(x-h)^2+k and y and x intercepts.

OpenStudy (campbell_st):

an easier way is o find the line of symmetry x = -b/2a in your question the line of symmetry is -2/(2 x -2) x = 1/2 then sub into the original equation to find the vertex y=-2 x(1/2)^2 + 2(1/2) - 3 y = -5/2 so the vertex is (1/2, -5/2) using in the required form y = -2(x - 1/2)^2 - 5/2 y intercept is when x = 0 so y intercept is y = -3 x intercepts, y = 0 or -2x^2 + 2x -3 = 0 the curve is negative definite so no x intercepts